please solve clearly in an organized way with a clear to read font. Follow the instructions and give a full answer.

2. Find the equation of the line tangent to the graph of x3 + y3 - 2xy = 0 at the point (1,1). (20 points)
Answer #1

Criven, find out the equation of the line tangent to the graph of x² + y²_quy zo at point (1,1) we know the equation of tangeequation of fargent, (2-11) dre y-y,= dy at nigg, equation of tangent at (11) y-! = dy (x-1) dre |(1,1) → 4-1 = -1 (1-1) ☆ y